OpenStudy (anonymous):

Find the sum of the first 12 terms of the sequence. 1, -4, -9, -14, . . . I have no idea how to approach this.

OpenStudy (amistre64):

is the sequence arithmetic, or geometric, or other?

OpenStudy (anonymous):

sequence arithmetic

OpenStudy (amistre64):

good, then it will have the general structure of:\[a_n=a_1+d(n-1)\]right?

OpenStudy (anonymous):

yes. I am just not sure how to work that structure into the given info.

OpenStudy (amistre64):

well, we are given a1,a2,a3, and a4 values, we just needed to determine "d" and we are pretty much set for the first part. so rearrange the general setup to solve for d\[\frac{a_n-a_1}{n-1}=d\] we know the value of a1=1 and the value of say, a2 = -4 sooo what do you determine the value of d is?

OpenStudy (anonymous):

How would i know 'n'?

OpenStudy (amistre64):

notice that when we n=2 the formula simply states that "d" is the difference between the first 2 numbers:\[\frac{a_2-a_1}{2-1}=a_2-a_1=d\]

OpenStudy (amistre64):

n is the "nth" number in the sequence

OpenStudy (anonymous):

so d =-5

OpenStudy (amistre64):

think of "n" as a placeholder. and yes, d=-5 :)

OpenStudy (amistre64):

soo\[a_n=a_1-5(n-1)~:~a_1=1\] \[a_n=1-5(n-1)\]now what would the 12th term be? let n=12

OpenStudy (anonymous):

a12=54 so would i just find each term up to the 12th and then add them all?

OpenStudy (amistre64):

well, -54 you can do that method if you want to. or you can work some simpler methods. they all give youthe same results

OpenStudy (anonymous):

ok, thank you so much! I think i'm beginning to understand this now :)

OpenStudy (amistre64):

youre welcome. let me know what your sum is, and i can dbl chk it if you want

OpenStudy (amistre64):

a_n = 1-5(n-1) = 1 -5n +5 = 6 - 5n a1 = 6 - 5(1) a2 = 6 - 5(2) a3 = 6 - 5(3) a4 = 6 - 5(4) ... a12 = 6 - 5(12) ---------------- sum: 12(6) - 5(1+2+3+4+5+...+12) 72 - 5(13)(6) 72 - 30(13) 72 - 390 is what i get

OpenStudy (anonymous):

so i got -318 with simply adding it, next time i will try your way. it seems quicker

OpenStudy (amistre64):

:) -318 is what i got too

OpenStudy (anonymous):

thanks again for your help!

OpenStudy (amistre64):

good luck ;)